Kyky232:

Type a polynomial with integer coefficients and a leading coefficient of 1. zeros-3, 0, 1 degree- 3

Vocaloid:

in general, if x = a is a zero, then x - a = 0 (where a is some arbitary value). also, if some product a*b*c = 0, then either a, b, or c must be 0 (more than one can be 0 at a time but at least one factor must be 0) so, applying this logic to our problem, if x = -3 is a zero, then (x - (-3)) = 0, or simplifying, (x+3) = 0 try to use this same logic for the other zeros. once you have your three expressions, you can simply multiply them together to get the polynomial. the leading coefficient will end up being 1, the coefficients will all be integers, and the degree will end up being 3, so all the conditions in the problem will end up being satisfied.